OpenStudy (anonymous):

Fair Tickets for 2 adults and 3 children cost $34. An adult ticket cost $2 more than a child ticket. what is the price of an adult ticket? Please Help !

OpenStudy (texaschic101):

2a + 3c = 34 c = a - 2 now just sub a - 2 in for c in the 1st equation and solve for a, the adult ticket

OpenStudy (anonymous):

so adult ticket costs $5.60 ?

OpenStudy (texaschic101):

2a + 3c = 34 2a + 3(a - 2) = 34 2a + 3a - 6 = 34 2a + 3a = 34 + 6 5a = 40 a = 8 c = a - 2 c = 8 - 2 c = 6 check... 2a + 3c = 34 2(8) + 3(6) = 34 16 + 18 = 34 34 = 34 (correct) Adult tickets cost $8.00 child's tickets cost $ 6.00

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

ohh I see my mistake I put a +6 instead of -6. thanks ! If you can help me on my next question that would be really helpful ;)